# Measurement to Meaning: A Validity-Centered Framework for AI Evaluation

### v1

@@ D.2 ImageNet
Table 6: An ImageNet (Deng et al., 2009; Russakovsky et al., 2014) Application. A subjective score for validity—the standard for “reasonable” is demonstrating that obvious risks to invalidity are addressed: [Uncaptioned image]: reasonable; [Uncaptioned image]: proceed with caution; [Uncaptioned image]: insufficient. Even for a score of “reasonable,” there will be weaknesses in the evidence. The score is given because the strengths outweigh the weaknesses in determining the validity of the claim from that evidence. This evaluation is an iterative process, acknowledging that both the benchmark and its interpretations may evolve over time.
@@
ImageNet (Deng et al., 2009; Russakovsky et al., 2014) (specifically ILSVRC 2012) is a benchmark for predicting an image’s label from a fixed set of 1000 diverse categories. The dataset—curated primarily from Flickr with human annotation—is evaluated using accuracy/error rate and precision/recall metrics.
@@
    • Strength: There is robust evidence that performance on ImageNet is both predictive of downstream task success (models excelling on ImageNet often perform well on benchmarks such as CIFAR or Caltech, and in real-world applications like wildlife classification (Norouzzadeh et al., 2018)) and concurrent with human-annotated labels under similar conditions (He et al., 2020; Zhai et al., 2022; Kornblith et al., 2018).

### v3

@@ D.2 ImageNet
Table 6: An ImageNet (Deng et al., 2009; Russakovsky et al., 2015) Application. A subjective score for validity—the standard for “reasonable” is demonstrating that obvious risks to invalidity are addressed: [Uncaptioned image]: reasonable; [Uncaptioned image]: proceed with caution; [Uncaptioned image]: insufficient. Even for a score of “reasonable,” there will be weaknesses in the evidence. The score is given because the strengths outweigh the weaknesses in determining the validity of the claim from that evidence. This evaluation is an iterative process, acknowledging that both the benchmark and its interpretations may evolve over time.
@@
ImageNet (Deng et al., 2009; Russakovsky et al., 2015) (specifically ILSVRC 2012) is a benchmark for predicting an image’s label from a fixed set of 1000 diverse categories. The dataset—curated primarily from Flickr with human annotation—is evaluated using accuracy/error rate and precision/recall metrics.
@@
    • Strength: There is robust evidence that performance on ImageNet is both predictive of downstream task success (models excelling on ImageNet often perform well on benchmarks such as CIFAR or Caltech, and in real-world applications like wildlife classification (Norouzzadeh et al., 2017)) and concurrent with human-annotated labels under similar conditions (He et al., 2020; Zhai et al., 2022; Kornblith et al., 2018).